It can be assumed that viral detection occurred with the same amount of DNA in S1 and S2 (S1 detected at cycle 35, S2 at cycle 32). How much more viral derived DNA was in the starting sample of S1 or S2? A. Twice as much B. Approximately 8 times more C. Approximately 16 times more D. Impossible to determine
Answer #1

Answer = B. Approximately 8 times more

S1 = 2^35 = 34359738368,   S2 = 2^32 = 4294967296
If S2 inital viral DNA double as S1 = 2X2^32 = 8589934592
If S2 inital viral DNA double as S1 = 8X2^32 = 34359738368